help me someone please!!
The length of a rectangle is 1 cm longer than its width. If the diagonal
of the rectangle is 4 cm, what are the dimensions (the length and the width) of the rectangle?

Answer by checkley75(3666)

X^2+(X+1)^2=4^2 X=WIDTH + (X+1)=LENGTH.
X^2+X^2+2X+1=16
2X^2+2X+1-16=0
2X^2+2X-15=0
using the quadratic equation we get:
x=(-2+-sqrt[2^2-4*2*-15])/2*2
x=(-2+-sqrt[4+120])/4
x=(-2+-sqrt124)/4
x=(-2+-11.1355)/4
x=(-2+11.1355)/4
x=9.1355/4
x=2.284 answer.
x=(-2-11.1355)/4
x=-13.1355/4
x=-3.2839 answer.